# FairKV: Balancing Per-Head KV Cache for Fast Multi-GPU Inference

### v1

@@ Abstract
KV cache techniques in Transformer models aim to reduce redundant computations at the expense of substantially increased memory usage, making KV cache compression an important and popular research topic. Recently, state-of-the-art KV cache compression methods implement imbalanced, per-head allocation algorithms that dynamically adjust the KV cache budget for each attention head, achieving excellent performance in single-GPU scenarios. However, we observe that such imbalanced compression leads to significant load imbalance when deploying multi-GPU inference, as some GPUs become overburdened while others remain underutilized. In this paper, we propose FairKV, a method designed to ensure fair memory usage among attention heads in systems employing imbalanced KV cache compression. The core technique of FairKV is Fair-Copying, which replicates a small subset of memory-intensive attention heads across GPUs using data parallelism to mitigate load imbalance. Theoretical analysis provides insights, while experiments on popular models, including LLaMA 70b and Mistral 24b model, demonstrate that FairKV increases throughput by 1.66x compared to standard tensor parallelism inference. Our code will be released as open source upon acceptance.
@@ 1.1 Background and Motivation
Traditional KV cache compression methods (StreamingLLMXiao et al. (2024), SnapKVLi et al. (2024) and PyramidKVCai et al. (2024)) naturally allocate an equal (fair) KV cache budget to all attention heads. This fair allocation simplifies the design and ensures uniform memory usage across heads. Recently, SOTA works of KV cache compression methods (AdaSnapKVFeng et al. (2024) and HeadKVFu et al. (2024)) introduce imbalanced per-head KV cache compression algorithms that dynamically adjust the KV cache budget for each attention head. These methods compress the KV cache more effectively and achieve SOTA performance in single-GPU scenarios.
@@ 2 Related Work
Table 2: GPU Utilization of Different Models. We applied Ada-SnapKV to three models, setting the KV cache budget to 128, 256, 512, and 1024, and measured its GPU utilization under tensor parallel sizes of 2, 4, and 8.
@@ 3.1 Observation of KV Cache Selection
The per-head KV cache compression algorithm results in varying KV cache budgets across different attention heads, leading to imbalanced workloads when combined with tensor parallelism. The results, summarized in Table 2, show a clear trend:
Decreasing GPU Utilization with Larger TP Sizes: For instance, in LLaMA-3.3-70B-Instruct with KV cache budget 128, GPU utilization drops from 92.5% (TP=2) to 81.6% (TP=4) and further to 64.7% (TP=8). A similar trend is observed in Meta-LLaMA-3-8B, where GPU utilization declines from 92.1% (TP=2) to 84.4% (TP=4) and 70.8% (TP=8) under the same KV cache budget.
Impact of KV cache Budget: Lower KV cache budgets generally yield better GPU utilization. For example, in Mistral-Small-24B-Instruct with TP=4, GPU utilization is 86.3% at KV=128 but decreases to 82.9% at KV=256 and further to 82.1% at KV=1024. Similarly, in LLaMA-3.3-70B-Instruct with TP=4, GPU utilization drops from 81.6% (KV=128) to 71.2% (KV=1024). This suggests that larger KV cache budgets lead to greater workload imbalance among attention heads, which negatively impacts GPU efficiency.
@@ 3.2 Empirical Performance Analysis
For the batch size latency model (Figure a), we observe that latency ($L$) increases approximately linearly with batch size ($B$) across different budget configurations. The relationship can be expressed as $L\approx\alpha B+\beta$, where $\alpha$ represents the slope and $\beta$ the initial offset. The graph shows four distinct budget levels (128, 256, 512, and 1024), with higher budget values corresponding to steeper slopes in the linear relationship. Similarly, the KV cache latency model (Figure b) demonstrates a comparable linear pattern, where latency ($L$) increases proportionally with the KV cache budget ($C$). This can be represented as $L\approx\gamma C+\delta$, where $\gamma$ and $\delta$ are the slope and offset parameters respectively. The data presents five different batch sizes (32, 64, 128, 256, and 512), with larger batch sizes showing more pronounced slopes in their linear relationships. In both cases, the approximately linear nature of these relationships is particularly noteworthy, as it suggests predictable scaling behavior in the system’s performance characteristics.
@@ 4.1 Overview
During inference with tensor parallelism, per-head KV cache compression algorithms can lead to imbalanced GPU workloads, which in turn reduces inference efficiency.To address this issue, we designed FairKV, a load-aware static approach that uses a search algorithm to reassemble and rearrange attention heads across layers, thereby balancing the load among GPUs. Moreover, FairKV leverages fair-copying mechanism to expand the search space via replication of attention heads and DataParallel techniques, enabling a more fine-grained balancing of GPU loads and enhancing GPU utilization during inference. Figure 2 visually illustrates the core concept of FairKV.
@@
Figure 2: Illustration of different head allocation strategies for multi-GPU inference in large-scale transformer models. The figure shows the following strategies: (1) Static Head Allocation (SHA), where attention heads are evenly distributed across GPUs without considering computational load; (2) Load-Aware Head Allocation (FairKV-NoDP), where attention heads are allocated based on their computational load, ensuring a balanced GPU utilization; (3) Load-Aware Head Allocation with DataParallel (FairKV-DP), where heads are replicated across GPUs for improved load distribution and efficiency;
@@ 4.2 Search Space
Fair-copying leverages Data Parallel techniques to expand the search space by replicating some redundant heads, with the goal of achieving more effective search outcomes. Experiments were conducted with selective head replication, and the results demonstrated improved latency distribution across GPUs. Allowing for limited redundancy provides an effective way to enhance parallel efficiency without excessive computational overhead.
4.3 Mathematical Model for Hybrid Parallelism in FairKV
4.3.1 System Parameters
  • $H_{l}=\{h_{1},...,h_{n}\}$: attention heads
  • $G=\{g_{1},...,g_{m}\}$: available GPUs
  • $w_{i}$: workload for head $h_{i}$
  • $r_{ij}$: replication factor of head $h_{i}$ on GPU $g_{j}$
4.3.2 Decision Variables
4.3.3 Parallelism Constraints
Data Parallelism Total replication factor for each head:
$\sum_{j\in G}r_{ij}\leq R_{max}\quad\forall i\in H_{l},\forall l\in L$ (3)
4.3.4 Optimization Objective
$\min\max_{j\in G}\sum_{l\in L}\sum_{i\in H_{l}}\frac{x_{ij}w_{i}}{r_{ij}}$ (4)
4.3.5 System Efficiency
@@ 4.4 Optimizer Design
We introduce a recursive backtracking algorithm that systematically explores possible head distributions while ensuring workload balance.
@@
Input: List of heads $L$, max GPU load $m$
Output: Optimal partitioning of $L$
Initialize result set $R$
procedure Backtrack($index,current$)
     if $index==|L|$ and $|current|\leq m$ then
         Add $current$ to $R$ return
     end if
     Add $L[index]$ to $current$ and recurse Backtrack($index+1,current$)
     Remove last element from $current$
     for $n=2$ to $m-|current|+1$ do
         Split $L[index]$ into $n$ parts and add to $current$ Backtrack($index+1,current$)
         Remove last $n$ elements from $current$
     end for
end procedure
This approach systematically searches for the best head distribution by iterating over all valid partitions. By ensuring balanced allocation, the algorithm minimizes inference latency across GPUs, leading to improved system efficiency.
@@ 4.5 Key Innovations
Latency-driven Optimization Backtracking minimizes inference latency, improving GPU utilization.
@@ 5.1 Experimental Setup
Model and Hardware Configurations We used the LLaMA-3.3-70B-Instruct model, the Meta-LLaMA-3-8B model, and the Mistral-Small-24B-Instruct-2501 model as our experimental models. Our hardware environment consists of four Nvidia A100-80G GPUs and 960GB of CPU memory.
@@
Baseline To the best of our knowledge, this study is the first to highlight the issue of GPU workload imbalance caused by per-head KV cache compression algorithms in tensor parallelism. Therefore, we chose the situation of conducting tensor parallel inference with Per-Head KV Compression but without FairKV as the baseline. By comparing the cases with and without FairKV, we can determine whether FairKV can improve GPU utilization and reduce inference latency.
Evaluation Metrics The purpose of the evaluation experiments is to verify the following two questions: First, whether FairKV can effectively reduce inference latency, and second, whether FairKV can effectively increase GPU utilization. Therefore, our evaluation metrics are the following two aspects: inference time and GPU utilization.
@@ 5.2.1 Throughput Improvement
We evaluated the performance of FairKV across different models. Specifically, we selected a diverse set of models, including LLaMA-3.3-70B-Instruct, Meta-LLaMA-3-8B, and Mistral-Small-24B-Instruct, and set the tensor parallel size to either 4 or 8 with RC fixed at 4. Throughput gains relative to SHA were measured under KV cache budgets of 128, 256, 512, and 1024. As shown in Figure 3, FairKV accelerates various models, and due to the inherent characteristics of each model, its throughput under SHA conditions differs. As a result, the acceleration effect of FairKV may vary. Among them, FairKV achieves the highest benefit of up to 1.66 on the Llama-3.3-70B-Instruct model.
@@
Figure 5: GPU Utilization with different Data Parallel Size on LLaMA-3.3-70B.
5.2.2 GPU Utilization Improvement
We used LLaMA-3.3-70B-Instruct to evaluate the impact of FairKV on GPU utilization. In Figure 4, we conducted ablation tests among standard model, FairKV without Fair-copying and FairKV with Fair-copying. The result indicate that both FairKV with or without Fair-copying significantly improves GPU utilization compared to standard model, demonstrating that the FairKV method can effectively balance GPU loads. Moreover, FairKV with Fair-copying shows further improvement over FairKV without Fair-copying. Then, to measure the impact of the parameter on the FairKV with Fair-copying group, we set the size of parallel size, which is also the count of copied heads (CH), to 1, 2, 3, and 4. We measured the GPU utilization for these groups under KV cache budgets of 128, 256, 512, and 1024, with the results shown in Figure 5. The results suggesting that incorporating only a small number of copied heads via Data Parallel can enhance the performance of the FairKV strategy greatly. We also observed a positive correlation between the performance of FairKV and CH; as CH increases, the GPU utilization curve becomes less steep, indicating that while increases in CH yield significant benefits when CH is small, the incremental gains diminish when CH is larger.
@@ 6 Conclusion
Our method has been extensively evaluated across various configurations, demonstrating consistent performance improvements. Specifically, FairKV significantly reduces inference latency and improves GPU utilization, regardless of the number of GPUs used, the model size (including different versions of LLaMa and Mistral), or the KV cache budget. This versatility highlights the robustness of our approach, making it effective across different settings and resource constraints. Experimental results show that FairKV achieves up to a 66% increase in inference throughput, with minimal impact on model accuracy. These findings confirm that FairKV is a promising solution for optimizing large-scale model inference in real-world, real-time applications.
Future work will explore the adaptability of FairKV to dynamic KV cache budgets and investigate its potential for scaling to models such as Mixture of Experts (MoE), where the computational load can vary significantly across different tasks. These enhancements aim to further extend the applicability of FairKV to a broader range of models and dynamic environments.
7 Limitations and Future works
Although FairKV provides significant improvements in inference efficiency, there are a few limitations that need to be addressed. First, FairKV is designed for a single machine with multiple GPUs and does not account for scenarios involving distributed systems across multiple machines. Second, the current method primarily focuses on the parallelization of the inference process without considering the separation of the prefill and decode stages, which is critical in certain real-time applications where the inference process is more complex and involves sequential decoding. Additionally, FairKV relies on the accuracy of the KV cache statistics for optimal head allocation. This means that the effectiveness of KV cache compression depends on the statistical properties of the cache, and any deviations from these properties could reduce the method’s performance. We will address these limitations in future work.
Appendix A Value of Per-Head KV Cache Compression Methods

### v2

@@ Abstract
KV cache techniques in Transformer models aim to reduce redundant computations at the expense of substantially increased memory usage, making KV cache compression an important and popular research topic. Recently, many popular KV cache compression methods implement imbalanced, per-head allocation algorithms that dynamically adjust the KV cache budget for each attention head, achieving excellent performance in single-GPU scenarios. However, we observe that such imbalanced compression leads to significant load imbalance when deploying multi-GPU inference, as some GPUs become overburdened while others remain underutilized. In this paper, we propose FairKV, a method designed to ensure fair memory usage among attention heads in systems employing imbalanced KV cache compression. The core technique of FairKV is Fair-Copying, which replicates a small subset of memory-intensive attention heads across GPUs using data parallelism to mitigate load imbalance. Theoretical analysis provides insights, while experiments on popular models, including LLaMA 70b and Mistral 24b model, demonstrate that FairKV increases throughput by 1.66x compared to standard tensor parallelism inference. Our code will be released as open source upon acceptance.
@@
11footnotetext: These authors contributed equally to this work.
@@ 1.1 Background and Motivation
Traditional KV cache compression methods (StreamingLLMXiao et al. (2024), SnapKVLi et al. (2024) and PyramidKVCai et al. (2024)) naturally allocate an equal (fair) KV cache budget to all attention heads. This fair allocation simplifies the design and ensures uniform memory usage across heads. Recently, many popular works of KV cache compression methods (AdaSnapKVFeng et al. (2024) and HeadKVFu et al. (2024)) introduce imbalanced per-head KV cache compression algorithms that dynamically adjust the KV cache budget for each attention head. These methods compress the KV cache more effectively and achieve impressive performance in single-GPU scenarios.
@@ 2 Related Work
Table 2: GPU Busy Rate of Different Models. We applied Ada-SnapKV to three models, setting the KV cache budget to 128, 256, 512, and 1024, and measured its GPU busy rate under tensor parallel sizes of 2, 4, and 8.
@@ 3.1 Observation of KV Cache Selection
The per-head KV cache compression algorithm results in varying KV cache budgets across different attention heads, leading to imbalanced workloads when combined with tensor parallelism. We introduce GPU busy rate, defined as the ratio of time the GPU spends actively executing inference computations to the total latency of the entire inference process, to quantify the resource underutilization caused by workload imbalance during parallel execution. The results, summarized in Table 2, show a clear trend:
Decreasing GPU Busy Rate with Larger TP Sizes: For instance, in LLaMA-3.3-70B-Instruct with KV cache budget 128, GPU busy rate drops from 92.5% (TP=2) to 81.6% (TP=4) and further to 64.7% (TP=8). A similar trend is observed in Meta-LLaMA-3-8B, where GPU utilization declines from 92.1% (TP=2) to 84.4% (TP=4) and 70.8% (TP=8) under the same KV cache budget.
Impact of KV Cache Budget: Lower KV cache budgets generally yield better GPU busy rate. For example, in Mistral-Small-24B-Instruct with TP=4, GPU busy rate is 86.3% at KV=128 but decreases to 82.9% at KV=256 and further to 82.1% at KV=1024. Similarly, in LLaMA-3.3-70B-Instruct with TP=4, GPU busy rate drops from 81.6% (KV=128) to 71.2% (KV=1024). This suggests that larger KV cache budgets lead to greater workload imbalance among attention heads, which negatively impacts GPU efficiency.
@@ 3.2 Empirical Performance Analysis
For the batch size latency model (Figure a), we observe that latency ($L$) increases approximately linearly with batch size ($B$) across different budget configurations. This is because during the decoding phase of large model inference, due to the autoregressive characteristic of Transformer-based architectures, the inference latency is primarily determined by GPU bandwidth rather than computational power (memory-bound). Consequently, as the batch size increases, the data transfer volume grows, making data transmission the dominant bottleneck, and latency increases nearly linearly. The relationship can be expressed as $L\approx\alpha B+\beta$, where $\alpha$ represents the slope and $\beta$ the initial offset. The graph shows four distinct budget levels (128, 256, 512, and 1024), with higher budget values corresponding to steeper slopes in the linear relationship.
Similarly, the KV cache latency model (Figure b) demonstrates a comparable linear pattern, where latency ($L$) increases proportionally with the KV cache budget ($C$). This occurs since the expanded KV Cache budget also amplifies data transfer demands, thereby maintaining the memory-bound bottleneck that dominates latency behavior. This can be represented as $L\approx\gamma C+\delta$, where $\gamma$ and $\delta$ are the slope and offset parameters respectively. The data presents five different batch sizes (32, 64, 128, 256, and 512), with larger batch sizes showing more pronounced slopes in their linear relationships.
In both cases, the approximately linear nature of these relationships is particularly noteworthy, as it suggests predictable scaling behavior in the system’s performance characteristics. Therefore, we conclude that during the decoding phase of model inference, neither increasing batch size nor expanding the KV Cache budget yields substantial efficiency improvements. This persistence of linear latency scaling fundamentally stems from the memory-bound nature of autoregressive generation in Transformer architectures, where GPU bandwidth constraints dominate computational throughput.
@@ 4.1 Overview
During inference with tensor parallelism, per-head KV cache compression algorithms can lead to imbalanced GPU workloads, which in turn reduces inference efficiency. To address this issue, we designed FairKV, a load-aware static approach that uses a search algorithm to reassemble and rearrange attention heads across layers, thereby balancing the load among GPUs. Moreover, FairKV leverages fair-copying mechanism to expand the search space via replication of attention heads and DataParallel techniques, enabling a more fine-grained balancing of GPU loads and enhancing GPU busy rate during inference. Figure 2 visually illustrates the core concept of FairKV.
@@
Figure 2: Illustration of different head allocation strategies for multi-GPU inference in large-scale transformer models. The figure shows the following strategies: (1) Static Head Allocation (SHA), where attention heads are evenly distributed across GPUs without considering computational load; (2) Load-Aware Head Allocation (FairKV-NoDP), where attention heads are allocated based on their computational load, ensuring a balanced GPU busy rate; (3) Load-Aware Head Allocation with DataParallel (FairKV-DP), where heads are replicated across GPUs for improved load distribution and efficiency;
@@ 4.2 Search Space
Fair-copying leverages Data Parallel techniques to expand the search space by replicating some redundant heads, with the goal of achieving more effective search outcomes. Experiments were conducted with selective head replication, and the results demonstrated improved latency distribution across GPUs. Allowing for limited redundancy provides an effective way to enhance parallel efficiency without excessive computational overhead. To better illustrate FairKV’s optimization objectives, we formulate a mathematical model, with detailed derivations provided in Appendix A.
@@ 4.3 Optimizer Design
We introduce a recursive backtracking algorithm that systematically explores possible head distributions and an optimal scheme selection algorithm to select the best attention head allocation scheme to promote workload balance.
We first use Algorithm 1 to enumerate all possible attention head replication schemes. Then, we apply a head-wise KV cache compression algorithm(like Ada-SnapKV) on a user-provided dataset to sample the KV cache allocation of each attention head per layer, which serves as the weight (denoted as W) for that layer. Finally, Algorithm 2 is employed for each layer to determine the optimal allocation scheme.
@@
1:List of heads $L$, maximum copied heads $m$
2:Result set $R$, all possible partitions of $L$
3:Initialize result set $R\leftarrow\emptyset$
[…]
5$\triangleright$ $\mathit{p}$ is a possible division of L
[…]
7:         $R\leftarrow R\cup\{\mathit{p}\}$
8:         return
9:     end if
[…]
11$\triangleright$ Case 1: Single copy
[…]
13:     Remove last element from $\mathit{p}$
[…]
16:         Append values of $L_{n}$ to $\mathit{p}$
17$\triangleright$ Case 2: Multiple copies
[…]
19:         Remove values of $L_{n}$ from $\mathit{p}$
20:     end for
21:end procedure
22:
23:Backtrack(00, $\emptyset$)
24$\triangleright$ Initial call with empty partition
25:return $R$
Algorithm 2 Optimal Scheme Selection
1:$R$ from Algorithm 1, attention heads weight $W$ of one layer, tensor parallelism size $tp$
2:best_scheme, optimal allocation scheme of one layer
[…]
5:for each $p$ in $R$ do
[…]
7$\triangleright$ Get occurrence counts
8:     $W_{A}\leftarrow\{W[i]/\!\text{copied\_size}[i]\ |\ \forall i\}$
9$\triangleright$ Adjust weights
10:     $\mathcal{P}\leftarrow\textsc{SplitIntoTpGroups}(p,tp)$
11$\triangleright$ Generate candidate schemes
12:     for each scheme in $\mathcal{P}$ do
13:         $\Delta\leftarrow\textsc{MaxGroupWeight}(W_{A})-\textsc{MinGroupWeight}(W_{A})$
[…]
17:         end if
18:     end for
19:end for
20:return best_scheme
Through these two algorithms, we can effectively improve GPU busy rates, thereby reducing idle waiting time across GPUs and enhancing inference efficiency.
@@ 4.4 Key Innovations
Latency-driven Optimization Backtracking minimizes inference latency, improving GPU busy rate.
@@ 5.1 Experimental Setup
Model and Hardware Configurations We used the LLaMA-3.3-70B-Instruct model, the Meta-LLaMA-3-8B model, and the Mistral-Small-24B-Instruct-2501 model as our experimental models. Our hardware environment consists of eightµ Nvidia A100-80G GPUs and 960GB of CPU memory.
@@
Baseline To the best of our knowledge, this study is the first to highlight the issue of GPU workload imbalance caused by per-head KV cache compression algorithms in tensor parallelism. Therefore, we chose the situation of conducting tensor parallel inference with Per-Head KV Compression but without FairKV as the baseline. By comparing the cases with and without FairKV, we can determine whether FairKV can improve GPU busy rate and reduce inference latency.
Evaluation Metrics The purpose of the evaluation experiments is to verify the following two questions: First, whether FairKV can effectively reduce inference latency, and second, whether FairKV can effectively increase GPU busy rate. Therefore, our evaluation metrics are the following two aspects: inference time and GPU busy rate.
@@ 5.2 Performance Evaluation
5.2.1 GPU busy rate Improvement
We used LLaMA-3.3-70B-Instruct to evaluate the impact of FairKV on GPU busy rate. In Figure 4, we conducted ablation tests among standard model, FairKV without Fair-copying and FairKV with Fair-copying. The result indicate that both FairKV with or without Fair-copying significantly improves GPU busy rate compared to standard model, demonstrating that the FairKV method can effectively balance GPU loads. Moreover, FairKV with Fair-copying shows further improvement over FairKV without Fair-copying. Then, to measure the impact of the parameter on the FairKV with Fair-copying group, we set the size of parallel size, which is also the count of copied heads (CH), to 1, 2, 3, and 4. We measured the GPU busy rate for these groups under KV cache budgets of 128, 256, 512, and 1024, with the results shown in Figure 5. The results suggesting that incorporating only a small number of copied heads via Data Parallel can enhance the performance of the FairKV strategy greatly. We also observed a positive correlation between the performance of FairKV and CH; as CH increases, the GPU busy rate curve becomes less steep, indicating that while increases in CH yield significant benefits when CH is small, the incremental gains diminish when CH is larger.
@@ 5.2.2 Throughput Improvement
We evaluated the performance of FairKV across different models. Specifically, we selected a diverse set of models, including LLaMA-3.3-70B-Instruct, Meta-LLaMA-3-8B, and Mistral-Small-24B-Instruct, and set the tensor parallel size to either 4 or 8 with CH fixed at 4. Throughput gains relative to SHA were measured under KV cache budgets of 128, 256, 512, and 1024. As shown in Figure 3, FairKV accelerates various models, and due to the inherent characteristics of each model, its throughput under SHA conditions differs. As a result, the acceleration effect of FairKV may vary. Among them, FairKV achieves the highest benefit of up to 1.66 on the Llama-3.3-70B-Instruct model.
@@
Figure 5: GPU Busy Rate with different Data Parallel Size on LLaMA-3.3-70B.
@@ 6 Conclusion
Our method has been extensively evaluated across various configurations, demonstrating consistent performance improvements. Specifically, FairKV significantly reduces inference latency and improves GPU busy rate, regardless of the number of GPUs used, the model size (including different versions of LLaMa and Mistral), or the KV cache budget. This versatility highlights the robustness of our approach, making it effective across different settings and resource constraints. Experimental results show that FairKV achieves up to a 66% increase in inference throughput, with minimal impact on model accuracy. These findings confirm that FairKV is a promising solution for optimizing large-scale model inference in real-world, real-time applications.
Future work will explore the adaptability of FairKV to dynamic KV cache budgets and investigate its potential for scaling to models such as Mixture of Experts (MoE), where the computational load can vary significantly across different tasks.
7 Limitations
Although FairKV provides significant improvements in inference efficiency, there are a few limitations that need to be addressed. First, FairKV is designed for a single machine with multiple GPUs and does not account for scenarios involving distributed systems across multiple machines. Second, the current method primarily focuses on the parallelization of the inference process without considering the separation of the prefill and decode stages, which is critical in certain real-time applications where the inference process is more complex and involves sequential decoding. Additionally, FairKV relies on the accuracy of the KV cache statistics for optimal head allocation. This means that the effectiveness of KV cache compression depends on the statistical properties of the cache, and any deviations from these properties could reduce the method’s performance.
Appendix A Mathematical Model for Hybrid Parallelism in FairKV
A.1 Mathematical Model for Hybrid Parallelism in FairKV
A.1.1 System Parameters
  • $H_{l}=\{h_{1},...,h_{n}\}$: attention heads
  • $G=\{g_{1},...,g_{m}\}$: available GPUs
  • $w_{i}$: workload for head $h_{i}$
  • $r_{ij}$: replication factor of head $h_{i}$ on GPU $g_{j}$
A.1.2 Decision Variables
A.1.3 Parallelism Constraints
Data Parallelism Total replication factor for each head:
$\sum_{j\in G}r_{ij}\leq R_{max}\quad\forall i\in H_{l},\forall l\in L$ (3)
A.1.4 Optimization Objective
$\min\max_{j\in G}\sum_{l\in L}\sum_{i\in H_{l}}\frac{x_{ij}w_{i}}{r_{ij}}$ (4)
A.1.5 System Efficiency
Appendix B Value of Per-Head KV Cache Compression Methods